OpenStudy (anonymous):

how to solve 2sintcost+sqrt(3)cos=0

OpenStudy (campbell_st):

well is the question \[2\sin(t)\cos(t) + \sqrt{3} \cos(t) = 0\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i believe the answer is 90,270,300,240 degrees

OpenStudy (campbell_st):

well what you need to do is factor out cos(t) so its \[\cos(t)(2\sin(t) + \sqrt{3})= 0\] so you are solving cos(t) = 0 and \[2\sin(t) - \sqrt{3} = 0\]